help on this please,,thanks
x+1= sqrt x+7

Answer by jai_kos(139)

x+1= sqrt x+7
Square on both sides, we get
(x+1)^2 = x +7
x^2 + x + 1 = x + 7
Get all the numbers on one side and “x” to the other side
x^2 + x – x = 7 -1
x^2 = 6
Take the sqrt of the equation, we get
x = sqrt(6)
Hence the value of “x” = sqrt(6)